# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v3

@@ Appendix B Author contributions
This project was a team effort, built on countless contributions from everyone involved. To acknowledge individual authors’ contributions and enable future inquiries to be directed appropriately, we followed the ACM’s policy on authorship (of Publications, 2024) and listed contributors for each part of the paper below.
B.1. Overall Author List and Contributions
@@
  • Hua Shen (University of Michigan, huashen@umich.edu): Initiated and led the overall project, prepared weekly project meetings, filtered papers, designed dimensions and codes (initial, revision), coded all papers, initiated the framework and developed human value and interaction modes analysis figures, participated in drafting all sections, paper revision and polishing.
@@
  • Tiffany Knearem (Google, tknearem@google.com): Led the HCI team, prepared weekly team meetings, filtered papers, designed dimensions and codes (initial, revision), coded partial papers, ideated the framework and analysis and future work content, participated in writing (Critical Thinking and AI Impact on Human sections), paper revision and polishing.
@@
  • Kundan Krishna (Carnegie Mellon University, kundank@andrew.cmu.edu): filtered papers, coded partial papers, ideated the framework and future work, participated in writing (Customizing AI section), designed dimensions and codes (initial, revision), paper revision and polishing, mainly involved in NLP Team

### v4

@@ 13 Author contributions
This project was a team effort, built on countless contributions from everyone involved. To acknowledge individual authors’ contributions and enable future inquiries to be directed appropriately, we followed the ACM’s policy on authorship acm_authorship_2024 and listed contributors for each part of the paper.
13.1 Overall Author List and Contributions
@@
  • Hua Shen (NYU Shanghai, New York University, huashen@nyu.edu): Initiated and led the overall project, prepared weekly project meetings, filtered papers, designed dimensions and codes (initial, revision), coded all papers, initiated the framework and developed human value and interaction modes analysis figures, participated in drafting all sections, paper revision and polishing.
@@
  • Tiffany Knearem (MBZUAI, Tknearem@gmail.com): Led the HCI team, prepared weekly team meetings, filtered papers, designed dimensions and codes (initial, revision), coded partial papers, ideated the framework and analysis and future work content, participated in writing (Critical Thinking and AI Impact on Human sections), paper revision and polishing.
@@
  • Kundan Krishna (Apple, kundank@andrew.cmu.edu): filtered papers, coded partial papers, ideated the framework and future work, participated in writing (Customizing AI section), designed dimensions and codes (initial, revision), paper revision and polishing, mainly involved in NLP Team